### v2

@@ 3 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol and its security under entangled attackers
Figure 4: Steps 2. and 3. of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol, where straight lines represent classical information and undulated lines represent quantum information. Image reproduced from [BCS22] with permission.
Consider a general attack to the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol, where Alice and Bob take the same role as in Section 2, but in addition, they have quantum registers that are able to hold $q$ qubits each, which can be possibly entangled. Since they will intercept the qubit sent by the $V_{0}$, the joint system will consist of $12q+$ 1 qubits. We now describe a general attack with attackers who pre-share entanglement.
Attack 3.2.
A general attack for a round of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol consists of (the superscripts denote dependence on $x$ and $y$, correspondingly):
@@
    Alice intercepts the qubit $Q$ and applies an arbitrary quantum operation to it and to her qubits, possibly entangling them. She keeps part of the resulting state, $q$ qubits at most, and sends the rest to Bob. Since the qubit $Q$ can be sent arbitrarily slow by $V_{0}$ (the verifiers only time the classical information), this happens before Alice and Bob can intercept $x$ and $y$. At this stage, Alice, Bob, and $V_{0}$ share a quantum state $\left|\psi\right\rangle$ of $12q+$ 1 qubits.
@@
    Alice and Bob intercept $x$ and $y$, and apply a unitary $U_{A}^{x}$ and $U_{B}^{y}$ on their local registers, respectively. Alice sends a part of her local state and $x$ to Bob and, Bob sends a part of his local state and $y$ to Alice. Denote by $\rho^{xy}$ their joint state.
@@
    Each party performs a POVM $\{A^{xy}_{a}\}$ and $\{B^{xy}_{b}\}$, $a,b\in\{0,1,\perp\}$, on their registers and Alice sends her outcome $a$ to $V_{0}$ and Bob sends his outcome $b$ to $V_{1}$.
A $q$-qubit strategy [BCS22] for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ consists of a starting state $\left|\psi\right\rangle$, corresponding to the starting state to attack the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol in step 1. before applying the unitary evolutions, unitaries $U_{A}^{x}$, $U_{B}^{y}$ and POVMs $\{A^{xy}_{a}\}$, $\{B^{xy}_{b}\}$, as described above. It can also be described by the tuple $\{\rho^{xy},A^{xy}_{a},B^{xy}_{b}\}_{x,y,a,b}$. The probabilities that the verifiers, after the attackers’ actions (for the random variable $\textsc{V}_{\textsc{AB}}$, as denoted above) record correct, wrong, no photon and different answers (the verifiers ‘Abort’ in such a case) are thus, respectively, given by
$\begin{split}\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]&=\frac{1}{2^{2n}}\sum_{a\in\{0,1\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\rho^{xy}V^{f(x,y)}_{a}\otimes A^{xy}_{a}\otimes B^{xy}_{a}\right],\\ \mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]&=\frac{1}{2^{2n}}\sum_{a\in\{0,1\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\rho^{xy}V^{f(x,y)}_{a}\otimes A^{xy}_{1-a}\otimes B^{xy}_{1-a}\right],\\ \mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{No photon}]&=\frac{1}{2^{2n}}\sum_{x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\rho^{xy}\mathbb{I}\otimes A^{xy}_{\perp}\otimes B^{xy}_{\perp}\right],\\ \mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}]&=\frac{1}{2^{2n}}\sum_{a\neq b\in\{0,1,\perp\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\rho^{xy}\mathbb{I}\otimes A^{xy}_{a}\otimes B^{xy}_{b}\right].\end{split}$ (24)
The criterion for a successful attack to a round of the protocol is the same as for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, i.e. (6) is fulfilled.
Our goal is to show that if the number of qubits that the attackers hold at the beginning of the protocol is linear, then given that they do not respond $\perp$’, their probability of being correct is strictly less than the corresponding probability of the honest prover. To this end, we define a relaxation of the condition of being correct, and we consider $q$-qubit strategies which have a high chance that the verifiers record Correct at the end of the protocol. More specifically, we will define a set of quantum states that are ‘good’ for a given fixed input, conditioned on actually playing. The first definition, which is an extension of Definition 4.1 in [BCS22], considers single round attacks that are ‘good’ for $l\leq 2^{2n}$ pars of $x,y$. The reason to do so is that the attackers could be wrong for pairs that might be asked with exponentially small probability.
@@
Let $\varepsilon\geq 0$ and $l\in\mathbb{N}$. A $q$-qubit strategy $\{\rho^{xy},A^{xy}_{a},B^{xy}_{b}\}_{x,y,a,b}$ for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ is $(\varepsilon$,l)-perfect if on $l$ pairs of strings $(x,y)$ if the attackers ‘respond’ with probability $\eta$ and
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq(1-p_{err})-\varepsilon.$ (25)
@@
Let $q\in\mathbb{N}$ be the number of qubits that each party Alice and Bob hold, and let $\varepsilon\geq 0$. We define $\mathcal{S}_{i}^{\varepsilon}$, for input $f(x,y)=i\in\{0,1\}$, as
@@
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}(i)\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq(1-p_{err})-\varepsilon,$ (27)
where $\textsc{Correct}(i)$ denotes being correct on input $i$, and
$\mathbb{P}[\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]=\eta.$ (28)
Notice that if $\left|\psi\right\rangle\in\mathcal{S}_{i}^{\varepsilon}$, then
$\begin{split}&\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}(i)\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq p_{err}+\varepsilon_{1},\\ &\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}(i)\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq\varepsilon_{2},\end{split}$ (29)
where $\varepsilon_{1},\varepsilon_{2}\geq 0$ are such that $\varepsilon_{1}+\varepsilon_{2}=\varepsilon$.
@@
where $(i)$ indicates ‘on the specific input $i$’, ‘$\mid\left|\phi\right\rangle$’ indicates that the probabilities we obtained with the state $\left|\phi\right\rangle$, $V_{a}^{i}$ are as in (9) and $\{A_{a}^{i},B_{a}^{i}\}_{a\in\{0,1,\perp\}}$ are POVMs. As a consequence of the SDP (19), considering that, from (13) and $\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]\leq p_{err}$, we have that $p_{win}+p_{err}\leq 1$, and thus we find that there exists a function $w:[0,1]\rightarrow[1-\cos^{2}(\frac{\pi}{8}),1]$ such that for all states $\left|\phi\right\rangle$, regardless of their dimension, upper bounds the performance of the attackers:
@@
Sketch (see below Lemma 3.10 for a detailed proof of the generalized version). We want to estimate the probability that for a randomly chosen $f$, we can find $f_{A}$ and $f_{B}$ such that the corresponding function $g$ is such that $\mathbb{P}_{x,y}[f(x,y)=g(f_{A}(x),f_{B}(y),\lambda)]\geq 3/4$. In a similar manner as in (76), we have that
$\mathbb{P}[f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\mathbb{P}_{x,y}[f(x,y)=g(f_{A}(x),f_{B}(y),\lambda)]]\leq 2^{(2^{n+1}+1)k}2^{2^{2n}h(1/4)}2^{-2^{2n}},$ (40)
@@
then
$\mathbb{P}[\textsc{V}_{\textsc{AB}}\neq\textsc{ Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\geq\frac{1}{4}[1-(\tilde{w}^{\xi}(\eta)+\Delta)].$ (42)
The proof of Lemma 3.11 is a particular case of the proof of Theorem 5.6. See Fig. 6 for a representation of the bound (42). As stated below, if $p_{err}$ is below these bounds, the attackers’ probability of being correct if they play is strictly smaller than the corresponding prover’s probability.
Theorem 3.12.
If the attackers respond with probability $\eta$ and control at most $q\leq\frac{n}{2}-5$ qubits at the beginning of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol and the honest prover’s error is such that $p_{err}<\frac{1}{4}[1-(\tilde{w}^{\xi}(\eta)+\Delta)]$, where $\Delta=0.013$, then the probability that the attackers are correct given that they respond is strictly smaller than the corresponding prover’s probability, i.e.
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{ Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]<\mathbb{P}[\textsc{V}_{\textsc{P}}=\textsc{ Correct}\mid\textsc{V}_{\textsc{P}}\neq\textsc{No photon}]=1-p_{err}.$ (43)
@@
Theorem 3.12 is an immediate consequence of Lemma 3.11. We see then that if $p_{err}<\frac{1}{4}[1-(\tilde{w}^{\xi}(\eta)+\Delta)]$ the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol is secure even in the presence of photon loss and attackers who pre-share entanglement. See Fig. 3 for the set of probabilities in $\Delta_{2}$ excluded by Theorem 3.12.

### v3

@@ 3 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol and its security under entangled attackers
[…]
$\cdot$
[…]
Figure 6: Schematic representation of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol, where straight lines represent classical information and undulated lines represent quantum information.
See Fig. 6 a schematic representation of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$. The $\mathrm{QPV}_{\mathrm{BB84}}^{f}$ protocol corresponds to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$  for $\eta=1$ and $p_{err}=0$. In the same way as for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, after $r$ rounds, the verifiers expect the answers to satisfy (1).
[…]
$T^{ent}_{i}(\textbf{a}_{i})=\gamma_{\textsc{c}}\textbf{1}_{\textsc{c}}(\textbf{a}_{i})-\gamma_{\perp}\textbf{1}_{\perp}(\textbf{a}_{i})-\gamma_{\textsc{i}}\textbf{1}_{\textsc{i}}(\textbf{a}_{i}),$ (30)
where $(\gamma_{\textsc{c}},\gamma_{\perp},\gamma_{\textsc{i}})=\frac{1}{\sqrt{488625947}}(943,1107,22057)$ and the superscript $ent$ is due to that this payoff will be used to securely verify the location even if attackers who pre-share entanglement intend to break $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$. Let
$\Gamma_{r}^{ent}=\sum_{i=1}^{r}T_{i}^{ent}(\textbf{a}_{i}),$ (31)
be the total score after $r$ rounds. We next introduce the following acceptance test with confidence level $\varepsilon_{\textrm{h}}$.
Definition 3.2.
Let $\varepsilon_{\textrm{h}}>0$. For the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol executed sequentially $r$ times, we define the acceptance test $\mathsf{T}^{rf-\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$, also referred to as the decision criterion, as follows: the verifiers accept the prover’s location if
$\Gamma_{r}^{ent}\geq r(\alpha^{ent}-\delta),$ (32)
where $\delta=\sqrt{\frac{4\ln(1/\varepsilon_{\textrm{h}})}{r}}$. Otherwise, they reject.
For an honest prover ($hp$), for every $i$,
$\mathbb{E}[T_{i}^{ent,hp}]=\gamma_{\textsc{c}}\eta(1-p_{err})-\gamma_{\perp}(1-\eta)-\gamma_{\textsc{i}}\eta p_{err}:\alpha^{ent}(\eta,p_{err}),$ (33)
and therefore, $\mathbb{E}[\Gamma_{r}^{ent,hp}]=r\alpha^{ent}$ for simplicity we will assume the dependence on $\eta$ and $p_{err}$ implicit. By Hoeffding’s inequality, see Lemma 2.3, with $T_{i}^{ent}\in[-\frac{22057}{\sqrt{488625947}},\frac{943}{\sqrt{488625947}}]\subset[-1,1]$, an honest prover will be accepted except with small probability at most $\varepsilon_{\textrm{h}}$, and therefore the test $\mathsf{T}^{rf-\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ is complete. We will show that, any attackers who pre-share a linear amount of qubits in $n$, after $r$ sequential executions of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, will fail the test with exponentially high probability, and thus showing soundness. Showing both completeness and soundness will prove that $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ is secure.
Similarly as $\mathsf{T}^{r\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$, this test is engineered from the analysis of the correlations attainable by attackers in Section 3, ensuring that they are rejected except with negligible probability; see the proof of Theorem 3.13 for details of its construction.
As introduced in Section 2, for the security analysis, we will consider the purified version of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, which is equivalent to it, and the difference relies on, $V_{0}$ preparing an EPR pair $\left|\Phi^{+}\right\rangle_{VP}$, keeping a qubit register $V$, and sending the register $P$ to the prover. In a later moment, the $V_{0}$ performs the measurement $\{V^{f(x,y)}_{v}=H^{f(x,y)}\left|v\rangle\langle v\right|_{V}H^{f(x,y)}\}_{v\in\{0,1\}}$ in his local register $V_{0}$.
Consider a general attack to the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol, where Alice and Bob take the same role as in Section 2, but in addition, prior to the execution of the protocol, they can pre-share entanglement. In the most general attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, Alice and Bob proceed as follows:
@@
    Alice intercepts the qubit state with register $P$ sent by $V_{0}$, and applies an arbitrary quantum operation to it and to a local register that she possesses, possibly entangling them. She keeps part of the resulting state, and sends the rest to Bob. Since the qubit $P$ can be sent arbitrarily slow by $V_{0}$ (the verifiers only time the classical information), this happens before Alice and Bob can intercept $x$ and $y$.
@@
    Alice intercepts $x$ and Bob intercepts $y$. At this stage, Alice, Bob, and $V_{0}$ share a quantum state $\left|\varphi\right\rangle$, make a partition and let $q$ be the number of qubits that Alice and Bob each hold, recall that $V_{0}$ holds a qubit with register $V$, and thus the three parties share a quantum state $\left|\varphi\right\rangle$ of $12q+$ 1 qubits. Alice and Bob apply a unitary $U_{A_{\text{k}}A_{\text{c}}}^{x}$ and $V_{B_{\text{k}}B_{\text{c}}}^{y}$ on their local registers $A_{\text{k}}A_{\text{c}}=:A$ and $B_{\text{k}}B_{\text{c}}=:B$, where k and c denote the registers that will be kept and communicated, respectively. Due to the Stinespring dilation, we consider unitary operations instead of quantum channels. They end up with the quantum state ${\left|\psi_{xy}\right\rangle=\mathbb{I}_{V}\otimes U_{A_{\text{k}}A_{\text{c}}}^{x}\otimes V_{B_{\text{k}}B_{\text{c}}}^{y}\left|\varphi\right\rangle}$.
@@
    Alice sends register $A_{\text{c}}$ and $x$ to Bob (and keeps register $A_{\text{k}}$), and Bob sends register $B_{\text{c}}$ and $y$ to Alice (and keeps register $B_{\text{k}})$.
  4. 4.
    Alice and Bob perform POVMs $\{A^{xy}_{a}\}_{a\in\{0,1,\perp\}}$ and $\{B^{xy}_{b}\}_{b\in\{0,1,\perp\}}$ on their local registers $A_{\text{k}}B_{\text{c}}=:A^{\prime}$ and $B_{\text{k}}A_{\text{c}}=:B^{\prime}$, and answer their outcomes $a$ and $b$ to their closest verifier, respectively.
[…]
Figure 7: Schematic representation of a general attack on $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$. Straight lines represent classical information, and undulated lines represent quantum information, including $x$ and $y$. The gray-shaded regions represent the corresponding tripartite quantum states in them.
[…]
$\begin{split}q_{\textsc{c}}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]&=\frac{1}{2^{2n}}\sum_{a\in\{0,1\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|V^{f(x,y)}_{a}\otimes A^{xy}_{a}\otimes B^{xy}_{a}\right],\\ q_{\textsc{w}}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\right]&=\frac{1}{2^{2n}}\sum_{a\in\{0,1\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|V^{f(x,y)}_{a}\otimes A^{xy}_{1-a}\otimes B^{xy}_{1-a}\right],\\ q_{\perp}\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\perp\right]&=\frac{1}{2^{2n}}\sum_{x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|\mathbb{I}_{V}\otimes A^{xy}_{\perp}\otimes B^{xy}_{\perp}\right],\\ q_{\lightning}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\right]&=\frac{1}{2^{2n}}\sum_{a\neq b\in\{0,1,\perp\},x,y\in\{0,1\}^{n}}\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|\mathbb{I}_{V}\otimes A^{xy}_{a}\otimes B^{xy}_{b}\right].\end{split}$ (34)
3.1 Security of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$
Our goal is to show that if the number of qubits $q$ that the attackers hold at the beginning of the protocol is linear, then, given that they do not respond $\perp$’, their probability of being correct is strictly less than the corresponding probability of the honest prover. To this end, we define a relaxation of the condition of being correct, and we consider $q$-qubit strategies which have a high chance that the verifiers record Correct at the end of the protocol. More specifically, we will define a set of quantum states that are ‘good’ for a given fixed input, conditioned on actually playing. The first definition, which is an extension of Definition 4.1 in [BCS22], considers single round attacks that are ‘good’ for $l\leq 2^{2n}$ pairs of $x,y$. The reason to do so is that the attackers could be wrong for pairs that might be asked with exponentially small probability.
@@ 3.1 Security of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$
Let $\varepsilon\geq 0$ and $l\in\mathbb{N}$. A $q$-qubit strategy
$\mathbf{S}_{\eta}^{f-BB84}=\{\left|\varphi\right\rangle,U^{x},V^{y},\{A^{xy}_{a}\}_{a},\{B^{xy}_{b}\}_{b}\}_{x,y}$
for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ is $(\varepsilon$,$l$)-perfect if there exists a set $\mathcal{L}$ with $\lvert\mathcal{L}\rvert\geq l$, such that the attackers
  1. 1.
    ‘respond’ with probability $\eta$ on these pairs:
    $\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|\mathbb{I}_{V}\otimes A^{xy}_{\perp}\otimes B^{xy}_{\perp}\right]=1-\eta\text{ }\forall(x,y)\in\mathcal{L},\$ (35)
  2. 2.
    and, up to $\varepsilon$, they are Correct at least with the same probability as an honest prover:
    $\sum_{a\in\{0,1\}}\mathrm{Tr}\left[\left|\psi_{xy}\rangle\langle\psi_{xy}\right|V^{f(x,y)}_{a}\otimes A^{xy}_{a}\otimes B^{xy}_{a}\right]\geq\eta\big{(}(1-p_{err})-\varepsilon\big{)}\text{ }\forall(x,y)\in\mathcal{L}.$ (36)
@@
Let Alice and Bob hold arbitrary registers $A$ and $B$, respectively, and let $\varepsilon\geq 0$. Let $V$ be a qubit register. For $i\in\{0,1\}$, we define the set $\mathcal{S}_{i}^{\varepsilon}$ as
@@
where
$\sum_{a\in\{0,1\}}\mathrm{Tr}\left[\left|\psi\rangle\langle\psi\right|V^{i}_{a}\otimes A^{i}_{a}\otimes B^{i}_{a}\right]\geq\eta\big{(}(1-p_{err})-\varepsilon\big{)},$ (38)
and
$\mathrm{Tr}\left[\left|\psi\rangle\langle\psi\right|\mathbb{I}_{V}\otimes A^{i}_{\perp}\otimes B^{i}_{\perp}\right]=1-\eta.$ (39)
@@
where $(i)$ indicates ‘on the specific input $i$’, ‘$\mid\left|\phi\right\rangle$’ indicates that the probabilities we obtained with the state $\left|\phi\right\rangle$, $V_{a}^{i}$ are as in (17) and $\{A_{a}^{i},B_{a}^{i}\}_{a\in\{0,1,\perp\}}$ are POVMs. As a consequence of the SDP (27), considering that, from (22) and $\Pr[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\leq\Pr[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]\leq p_{err}$, we have that $p_{win}+p_{err}\leq 1$, and thus we find that there exists a function $w:[0,1]\rightarrow[1-\cos^{2}(\frac{\pi}{8}),1]$ such that for all states $\left|\phi\right\rangle$, regardless of their dimension, upper bounds the performance of the attackers:
@@
Sketch (see below Lemma 3.10 for a detailed proof of the generalized version). We want to estimate the probability that for a randomly chosen $f$, we can find $f_{A}$ and $f_{B}$ such that the corresponding function $g$ is such that $\Pr_{x,y}[f(x,y)=g(f_{A}(x),f_{B}(y),\lambda)]\geq 3/4$. In a similar manner as in (84), we have that
$\Pr[f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\Pr_{x,y}[f(x,y)=g(f_{A}(x),f_{B}(y),\lambda)]]\leq 2^{(2^{n+1}+1)k}2^{2^{2n}h(1/4)}2^{-2^{2n}},$ (50)
@@
then, for any $q$-qubit strategy for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, the probability that the attackers answer Correct is at most
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]\leq\eta\big{(}1-\frac{1}{4}[1-(\tilde{w}^{\epsilon}(\eta)+\Delta)]\big{)}.$ (52)
The proof of Theorem 3.11 is a particular case of the proof of Theorem 5.6. See Fig. 9 for a representation of the bound (52) and see Fig. 5 for the set of probabilities in $\Delta_{2}$ excluded by this theorem. Notice that if $p_{err}<1-\frac{1}{4}[1-(\tilde{w}^{\epsilon}(\eta)+\Delta)]$, then the probability that the attackers’ answer Correct is stricly below the corresponding probability for an honest prover.